# DeepLag: Discovering Deep Lagrangian Dynamics for Intuitive Fluid Prediction

### v3

@@ Abstract
Accurately predicting the future fluid is vital to extensive areas such as meteorology, oceanology, and aerodynamics. However, since the fluid is usually observed from the Eulerian perspective, its moving and intricate dynamics are seriously obscured and confounded in static grids, bringing thorny challenges to the prediction. This paper introduces a new Lagrangian-Eulerian combined paradigm to tackle the tanglesome fluid dynamics. Instead of solely predicting the future based on Eulerian observations, we propose DeepLag to discover hidden Lagrangian dynamics within the fluid by tracking the movements of adaptively sampled key particles. Further, DeepLag presents a new paradigm for fluid prediction, where the Lagrangian movement of the tracked particles is inferred from Eulerian observations, and their accumulated Lagrangian dynamics information is incorporated into global Eulerian evolving features to guide future prediction respectively. Tracking key particles not only provides a transparent and interpretable clue for fluid dynamics but also makes our model free from modeling complex correlations among massive grids for better efficiency. Experimentally, DeepLag excels in three challenging fluid prediction tasks covering 2D and 3D, simulated and real-world fluids.
@@ 1 Introduction
Fluids, characterized by a molecular structure devoid of resistance to external shear forces, readily undergo deformation even under minimal force, thereby typically presenting extremely complex dynamics [10]. Consequently, the solvability of theorems for fluid mechanics, such as Navier-Stokes equations, is confined to a limited subset of flows due to the complex form and intricate multiphysics interactions [38]. In practical applications, computational fluid dynamics (CFD) is widely used to infer future fluid by numerical simulations. However, it suffers from substantial computation costs due to the curse of dimensionality [40]. Accurately predicting future dynamics is still a challenging problem. Recently, benefiting from the tremendous non-linear modeling capability, deep models [8, 28, 22] have emerged as a promising tool for fluid prediction, which can be trained with CFD simulations or real-world observations and perform as a fast surrogate model during inference.
Figure 1: Comparison between Lagrangian and Eulerian perspectives. The left part presents the trajectory of five representative particles learned by DeepLag on the Ocean Current dataset, which provides a Lagrangian perspective for Eulerian-recorded fluid on the right.
@@
In parallel to the Eulerian method, we notice another major approach for elucidating fluid dynamics, the Lagrangian method [13], also known as the tracking method. This method primarily focuses on tracing individual fluid particles by conducting the temporal evolution of the position and velocity of each particle. Unlike the Eulerian methods, the Lagrangian approach describes the fluid dynamics through the moving trajectory of individual fluid particles, offering a more natural and neat representation of fluid dynamics with inherent advantages in capturing intricate flow dynamics. Moreover, it helps get around the Courant–Friedrichs–Lewy condition [6] for advection, allowing a larger inference step while retaining stability. In Figure 1, we can find that fluid dynamics is much more visually apparent in Lagrangian trajectories than density changes on static Eulerian grids.
Building on the two perspectives mentioned earlier, we propose DeepLag as an Eulerian-Lagrangian Recurrent Network. Our aim is to integrate Lagrangian tracking into the deep model, thereby enhancing the dynamics modeling in Eulerian fluid prediction. To achieve this, we present the EuLag Block, a powerful module that accomplishes Lagrangian tracking and Eulerian predicting at various scales. By leveraging the cross-attention mechanism, the EuLag Block assimilates tracked Lagrangian particle dynamics into the Eulerian field, guiding fluid prediction. It also forecasts the trajectory and dynamics of Lagrangian particles with the aid of Eulerian features. This unique Eulerian-Lagrangian design harnesses the dynamics information captured by Lagrangian trajectories and the fluid-structure features learned in the Eulerian grid. In our experiments, DeepLag consistently outperforms existing models, demonstrating state-of-the-art performance across three representative datasets, covering 2D and 3D fluid at various scales. Our contributions are as follows:
  • Going beyond learning fluid dynamics at static grids, we propose DeepLag featuring the Eulerian-Lagrangian Recurrent Network, which is the first deep fluid prediction model that explicitly combines Eulerian and Lagrangian frameworks from fluid dynamics concisely.
@@ 2.2 Neural Fluid Prediction
ODE-based generative models
As depicted in Figure 2(a), these methods [4, 26, 55] encode multivariate fields into a single-variable latent state $\mathbf{z}$, on which they model an ODE governing the state function $\mathbf{z}_{t}:\mathcal{T}\rightarrow\mathbb{R}^{d}$, representing the first-order time derivative, through neural networks. However, the absence of physical meaning and guidance for latent states in evolution leads to error accumulation and a generally short forecasting horizon. A detailed comparison between DeepLag and these models is provided in Appendix B.
@@ 3 DeepLag
Following the convention [21], we formalize the fluid prediction problem as learning future fluid given past observation, as shown in Figure 2(d). Given a bounded open subset of $d$-dimensional Euclidean space $\mathcal{D}\subset\mathbb{R}^{d}$ and an Eulerian space $\mathcal{U}\subset\mathbb{R}^{o}$ with $o$ observed physical quantities, $\mathbf{u}_{t}(\mathbf{x})\subset\mathcal{U}$ and $\mathbf{u}_{t+1}(\mathbf{x})\subset\mathcal{U}$ represent the Eulerian fluid field observation at two consecutive time steps on a finite coordinate set $\mathbf{x}\subset\mathcal{D}$. Provided $P$ step observations ${U}_{P}=\{{\mathbf{u}}_{1},\ldots,{\mathbf{u}}_{P}\}$, the fluid prediction process can be written as the following autoregressive paradigm:
@@ 3.1 Overall Framework
To better capture the complex dynamics within the fluid field, we sample crucial points from the given observations in the Eulerian perspective to determine the initial position for Lagrangian tracking in the dynamic sampling module. Explicitly, for the first predicting step, i.e. $t=P$, given observations ${\mathbf{u}}_{t}=\{{\mathbf{u}}_{t}(\mathbf{x}_{k})|\mathbf{x}_{k}\in\mathcal{D}_{l},1\leq k\leq N_{l}\}\in\mathbb{R}^{N_{l}\times o}$ on all $N_{l}$ points in the observation domain $\mathcal{D}_{l}\subset\mathbb{R}^{d}$ of each scale, we extract their spatial dynamics by a convolutional network. We then calculate the probability matrix $\mathbf{S}\in\mathbb{R}^{N_{l}}$ using softmax along the spatial dimension:
@@
where $\operatorname{ConvNet}(\cdot)$ consists of a convolutional, activation, and fully-connected layer. We then sample $M_{l}$ Lagrangian tracking particles using the probability matrix at each scale,
@@
where the scale index $l$ is omitted. The EuLag Block optimally leverages the complementary strengths of Eulerian and Lagrangian representations, facilitating enhanced training and mutual refinement between these two perspectives. More details about the implementation of the EuLag Block are expanded in the following subsection.
@@ 3.2 EuLag Block
Figure 3: Overview of the EuLag Block, which accumulates previous dynamics information to guide Eulerian evolution for predicting particle movement. Scale index $l$ is omitted.
@@
where $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$ and $\mathbf{W}_{V}$ stand for linear projection layers. We wrap the above attention into a standard transformer block with residual as Lagrangian-dynamics-guided Eulerian Evolution process: $\hat{\mathbf{u}}_{t+1}={\mathbf{u}}_{t}+\operatorname{LagToEu}({\mathbf{u}}_{t},{\mathbf{h}}_{t}^{\prime},{\mathbf{h}}_{t}^{\prime})$.
@@
where $\operatorname{TrmBlock}_{Q,K,V}(\cdot,\cdot)$ represents standard transformer block with $\mathbf{W}^{{}^{\prime}}_{Q},\mathbf{W}^{{}^{\prime}}_{K},\mathbf{W}^{{}^{\prime}}_{V}$ as linear projection matrix. The forecasted global Lagrangian dynamic and movement of tracking particles are produced by two transformers with the same query and key projection but different value projections. To better model the dynamic evolution of the particles, we gather local Lagrangian dynamic by employing bilinear interpolation to output Eulerian features $\tilde{\mathbf{u}}_{t+1}$ on particles $\mathbf{p}_{t+1}$, then use a linear function to interact with global dynamics information ${\mathbf{h}}_{\text{global},t+1}$, naming ${\mathbf{h}}_{t+1}=\operatorname{Aggregate}(\tilde{\mathbf{u}}_{t+1},\mathbf{p}_{t+1},{\mathbf{h}}_{\text{global},t+1})$. Additionally, as the input flow field may have an open boundary, the particles could move out of the observation domain. We check the position of the particles and resample them in case they exit. Similar to the previous Eulerian Evolution process, the above Lagrangian Tracking process is summarized as $\operatorname{EuToLag}(\cdot)$.
@@ 4 Experiments
We evaluated DeepLag on three challenging benchmarks, including simulated and real-world scenarios, covering both 2D and 3D, as well as single and multi-physics fluids. Following the previous convention [21], we train DeepLag and the baselines for each task to predict ten future timesteps in an autoregressive fashion given ten past observations. Detailed benchmark information is listed in Table 1. We provide an elaborate analysis of the efficiency, parameter count, and performance difference in section 4.4 and Appendix E. Additionally, more detailed visualizations, besides in later this section, are provided in Appendix H. Furthermore, the trained models are engaged to perform 30 frames extrapolation to examine their long-term stability.
@@ 10 Frames 30 Frames
Figure 4: Showcases and timewise relative L2 on Bounded Navier-Stokes dataset. Both predictions and absolute error maps are plotted.
@@
To intuitively present the forecasting skills of different models, we also provide showcase comparisons in Figure 4. We can find that DeepLag can precisely illustrate the vortex in the center of the figure and the Kármán vortex phenomenon formed behind the upper left pillar. As for U-Net and LSM, although they successfully predicted the position of the center vortex, the error map shows that they failed to predict the density field accurately. In addition, FactFormer deteriorates on this benchmark. This may be because it is based on spatial factorization, which is unsuitable for irregularly placed boundary conditions. These results further highlight the benefits of Eulerian-Lagrangian co-design, which can simultaneously help with dynamic and density prediction.
@@ 10 Days 30 Days
Setups
Predicting large-scale ocean currents, especially in regions near tectonic plate boundaries prone to disasters such as tsunamis due to intense terrestrial activities, plays a crucial role in various domains. Hence, we also explore this valuable scenario in our experiments. More details about this benchmark can be found in Appendix C.2.
Quantitive results
@@
We presented different model predictions in Figure 5. In comparison to other models, DeepLag exhibits the most minor prediction error. It accurately predicts the location of the high-temperature region to the south area and provides a clear depiction of the Kuroshio pattern [37] bounded by the red box.
@@
To reveal the effect of learning Lagrangian trajectories, we visualize tracked particles in Figure 5. We observe that all the particles move from west to east, consistent with the Pacific circulation. Additionally, the tracked particles show distinct moving patterns, confirming their ability to represent complex dynamics. The movement of upper particles matches the sinuous trajectory of the Kuroshio current, demonstrating the capability of DeepLag to provide interpretable evidence for prediction results. Visualizing the tracking points in Lagrangian space instills confidence in the reliability and interpretability of the predictions made by our system.
@@
Table 4: Performance comparison on the 3D Smoke dataset. The relative L2 with promotion is recorded.
@@
3D fluid prediction has been a long-standing challenge due to the tanglesome dynamics involved. Therefore, we created this benchmark to describe a scenario in which smoke flows under the influence of buoyancy in a three-dimensional bounding box. For more details, please refer to Appendix C.3.
@@
In Figure 6, we compared prediction results on the 3D Smoke dataset. DeepLag performs well in capturing the convection and diffusion of smoke within the bounding box. In comparison, the predictions of U-Net tend to predict the averaged value across various surfaces, resulting in blurred details, which also indicates its deficiency in dynamics modeling. As for LSM and FactFormer, we can find that they present more errors in the smoke boundaries that usually involve complex waves. In contrast, our model excels in overall and cross-section error and performs excellently in predicting subtle flows.
Figure 6: Showcases comparison of the whole space (left part) and a cross-section (right part) on 3D Smoke dataset. For better visualization, we present the absolute value of the prediction error and normalize the whole space error into $(0,0.12)$. As for the cross-section visualization, we choose the $xOy$ plane in the middle 3D fluid and normalize error maps to $(-0.5,0.5)$.

### v4

@@ Abstract
Accurately predicting the future fluid is vital to extensive areas such as meteorology, oceanology, and aerodynamics. However, since the fluid is usually observed from the Eulerian perspective, its moving and intricate dynamics are seriously obscured and confounded in static grids, bringing thorny challenges to the prediction. This paper introduces a new Lagrangian-Eulerian combined paradigm to tackle the tanglesome fluid dynamics. Instead of solely predicting the future based on Eulerian observations, we propose DeepLag to discover hidden Lagrangian dynamics within the fluid by tracking the movements of adaptively sampled key particles. Further, DeepLag presents a new paradigm for fluid prediction, where the Lagrangian movement of the tracked particles is inferred from Eulerian observations, and their accumulated Lagrangian dynamics information is incorporated into global Eulerian evolving features to guide future prediction respectively. Tracking key particles not only provides a transparent and interpretable clue for fluid dynamics but also makes our model free from modeling complex correlations among massive grids for better efficiency. Experimentally, DeepLag excels in three challenging fluid prediction tasks covering 2D and 3D, simulated and real-world fluids. Code is available at this repository: https://github.com/thuml/DeepLag.
@@ 1 Introduction
Fluids, characterized by a molecular structure that offers no resistance to external shear forces, easily deform even under minimal stress, leading to highly complex and often chaotic dynamics [10]. Consequently, the solvability of fundamental theorems in fluid mechanics, such as Navier-Stokes equations, is constrained to only a limited subset of flows due to their inherent complexity and intricate multiphysics interactions [38]. In practical applications, computational fluid dynamics (CFD) is widely employed to predict fluid behavior through numerical simulations, but it is hindered by significant computational costs. Accurately forecasting future fluid dynamics remains a formidable challenge. Recently, deep learning models [8, 28, 22] have shown great promise for fluid prediction due to their exceptional non-linear modeling capabilities. These models, trained on CFD simulations or real-world data, can serve as efficient surrogate models, dramatically accelerating inference.
Figure 1: Comparison between Lagrangian (left) and Eulerian (right) perspectives. The left depicts the learned trajectories of Lagrangian particles overlaid on the mean state, while the right displays the positions of tracked particles in successive Eulerian frames. Fluid motion is more visibly represented through the dynamic Lagrangian view compared to the density variations in static Eulerian grids.
@@
In parallel to the Eulerian method, we notice another major approach for elucidating fluid dynamics, the Lagrangian method [13], also known as the particle tracking method. This method primarily focuses on tracing individual fluid particles by modeling the temporal evolution of their position and velocity. Unlike the Eulerian methods, which observe fluid flow at fixed spatial locations, the Lagrangian approach describes the fluid dynamics through the moving trajectory of individual fluid particles, offering a more natural and neat representation of fluid dynamics with inherent advantages in capturing intricate flow dynamics. Moreover, it allows a larger inference time step than Eulerian methods while complying with the Courant–Friedrichs–Lewy condition [6] that guarantees stability. In Figure 1, we can find that fluid dynamics is much more visually apparent in Lagrangian trajectories on the left compared to the density changes observed on static Eulerian grids on the right.
Building on the two perspectives mentioned earlier, we propose DeepLag as a Eulerian-Lagrangian Recurrent Network. Our aim is to integrate Lagrangian tracking into the deep model, thereby enhancing the dynamics modeling in Eulerian fluid prediction. To achieve this, we present the EuLag Block, a powerful module that accomplishes Lagrangian tracking and Eulerian predicting at various scales. By leveraging the cross-attention mechanism, the EuLag Block assimilates tracked Lagrangian particle dynamics into the Eulerian field, guiding fluid prediction. It also forecasts the trajectory and dynamics of Lagrangian particles with the aid of Eulerian features. This unique Eulerian-Lagrangian design harnesses the dynamics information captured by Lagrangian trajectories and the fluid-structure features learned in the Eulerian grid. In our experiments, DeepLag consistently outperforms existing models, demonstrating state-of-the-art performance across three representative datasets, covering 2D and 3D fluid at various scales. Our contributions are as follows:
  • Going beyond learning fluid dynamics at static grids, we propose DeepLag featuring the Eulerian-Lagrangian Recurrent Network, which integrates both Eulerian and Lagrangian frameworks from fluid dynamics within a pure deep learning framework concisely.
@@ 2.2 Neural Fluid Prediction
Classical ML methods
As depicted in Figure 2(a), these methods either replace part of a numerical method with a neural surrogate [39] or encode multivariate fields into a single-variable latent state $\mathbf{z}$ [4, 26, 55], on which they model an ODE governing the state function $\mathbf{z}_{t}:\mathcal{T}\rightarrow\mathbb{R}^{d}$, representing the first-order time derivative, through neural networks. However, the absence of physical meaning and guidance for latent states in evolution leads to error accumulation and a generally short forecasting horizon. A detailed comparison between DeepLag and these models is provided in Appendix B.
@@ 3 DeepLag
Following the convention of neural fluid prediction [21], we formalize the fluid prediction problem as learning future fluid given past observation, as shown in Figure 2(d). Given a bounded open subset of $d$-dimensional Euclidean space $\mathcal{D}\subset\mathbb{R}^{d}$ and an Eulerian space $\mathcal{U}\subset\mathbb{R}^{o}$ with $o$ observed physical quantities, let $\mathbf{u}_{t}(\mathbf{x})\subset\mathcal{U}$ and $\mathbf{u}_{t+1}(\mathbf{x})\subset\mathcal{U}$ represent the Eulerian fluid field observation at two consecutive time steps on a finite coordinate set $\mathbf{x}\subset\mathcal{D}$, we aim at fitting the mapping $\Phi:\mathbf{u}_{t}(\mathbf{x})\rightarrow\mathbf{u}_{t+1}(\mathbf{x})$. Concretely, provided initial $P$ step observations ${U}_{P}=\{{\mathbf{u}}_{1},\ldots,{\mathbf{u}}_{P}\}$, the fluid prediction process can be written as the following autoregressive paradigm:
@@ 3.1 Overall Framework
To better capture the complex dynamics within the fluid field, we sample crucial points from the observations in the Eulerian perspective to determine the initial position for Lagrangian tracking in the dynamic sampling module. Explicitly, for the first predicting step $t=P$, given observations ${\mathbf{u}}_{t}=\{{\mathbf{u}}_{t}(\mathbf{x}_{k})|\mathbf{x}_{k}\in\mathcal{D}_{l},1\leq k\leq N_{l}\}\in\mathbb{R}^{N_{l}\times o}$ on all $N_{l}$ points in the observation domain $\mathcal{D}_{l}\subset\mathbb{R}^{d}$ of each scale, we extract their spatial dynamics by a convolutional network. We then calculate the probability matrix $\mathbf{S}\in\mathbb{R}^{N_{l}}$ via softmax along the spatial dimension:
@@
where $\operatorname{ConvNet}(\cdot)$ consists of a convolutional, activation, and fully-connected layer. We then sample $M_{l}$ Lagrangian tracking particles using the probability matrix at each scale:
@@
where the scale index $l$ is omitted for simplicity and clarity in notation. The EuLag Block optimally leverages the complementary strengths of Eulerian and Lagrangian representations, facilitating enhanced training and mutual refinement between these two perspectives. More details about the specific implementation of the EuLag Block are expanded upon in the following subsection 3.2.
@@ 3.2 EuLag Block
Figure 3: Overview of the EuLag Block, which accumulates previous dynamics information to guide Eulerian evolution for predicting particle movement. Scale index $l$ is omitted for simplicity.
@@
where $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$ and $\mathbf{W}_{V}$ stand for linear projection layers. We wrap the above attention into a standard transformer block with residual as Lagrangian-dynamics-guided Eulerian Evolution process:
$\begin{split}\displaystyle\tilde{\mathbf{u}}_{t+1}={\mathbf{u}}_{t}+\operatorname{LagToEu}({\mathbf{u}}_{t},{\mathbf{h}}_{t}^{\prime},{\mathbf{h}}_{t}^{\prime}).\end{split}$ (8)
@@
where $\operatorname{TrmBlock}_{Q,K,V}(\cdot,\cdot)$ represents standard transformer block with $\mathbf{W}^{{}^{\prime}}_{Q},\mathbf{W}^{{}^{\prime}}_{K},\mathbf{W}^{{}^{\prime}}_{V}$ as linear projection matrix. The forecasted global Lagrangian dynamic and movement of tracking particles are produced by two transformers with the same query and key projection but different value projections. To better model the dynamic evolution of the particles, we gather local Lagrangian dynamic by employing bilinear interpolation to output Eulerian features $\tilde{\mathbf{u}}_{t+1}$ on new particle position $\mathbf{p}_{t+1}$, then use a linear function to interact it with global dynamics information ${\mathbf{h}}_{\text{global},t+1}$, naming ${\mathbf{h}}_{t+1}=\operatorname{Aggregate}(\tilde{\mathbf{u}}_{t+1},\mathbf{p}_{t+1},{\mathbf{h}}_{\text{global},t+1})$. Additionally, as the input flow field may have an open boundary, the particles could move out of the observation domain. We check the updated position of the particles and resample the ones that exit from the latest probability matrix $\mathbf{S}$. Similar to the previous Eulerian Evolution process, the above Lagrangian Tracking process can be summarized as:
$\begin{split}\displaystyle\mathbf{h}_{t+1},\mathbf{p}_{t+1}=\operatorname{EuToLag}({\mathbf{h}}_{t}^{\prime},\tilde{\mathbf{u}}_{t+1},\tilde{\mathbf{u}}_{t+1}).\end{split}$ (10)
@@ 4 Experiments
We evaluated DeepLag on three challenging benchmarks, including simulated and real-world scenarios, covering both 2D and 3D, as well as single and multi-physics fluids. Following the previous convention [21], we train DeepLag and the baselines for each task to predict ten future timesteps in an autoregressive fashion given ten past observations. Detailed benchmark information is listed in Table 1. We provide an elaborate analysis of the efficiency, parameter count, and performance difference in section 4.4 and Appendix E. Additionally, more detailed visualizations, besides in later this section, are provided in Appendix H. Furthermore, the trained models are engaged to perform 100 frames extrapolation to examine their long-term stability, whose results are in Appendix I.
@@ 10 Frames 30 Frames
Figure 4: Showcases (left) and timewise relative L2 (right) on Bounded Navier-Stokes dataset. Both predictions (upper row) and absolute error maps (lower row) are plotted for intuitive comparison.
@@
To intuitively present the forecasting skills of different models, we also provide showcase comparisons in Figure 4 and particle movements predicted by DeepLag in Appendix K. We can find that DeepLag can precisely illustrate the vortex in the center of the figure and give a reasonable motion mode of the Kármán vortex phenomenon formed behind the upper left pillar. As for U-Net and LSM, although they successfully predicted the position of the center vortex, the error map shows that they failed to predict the density field as accurately as DeepLag. In addition, FactFormer deteriorates on this benchmark. This may be because it is based on spatial factorization, which is unsuitable for irregularly placed boundary conditions. These results further highlight the benefits of Eulerian-Lagrangian co-design, which can simultaneously help with dynamic and density prediction.
@@ 4.2 Ocean Current
Setups
Predicting large-scale ocean currents, especially in regions near tectonic plate boundaries prone to disasters such as tsunamis due to intense terrestrial activities, plays a crucial role in various domains. Hence, we also explore this challenging real-world scenario in our experiments. More details about the source and settings of this benchmark can be found in Appendix C.2.
Quantitive results
@@ 10 Days 30 Days
To provide an intuitive comparison, we plotted different model predictions in Figure 5. In comparison to other models, DeepLag exhibits the most minor prediction error. It accurately predicts the location of the high-temperature region to the south area and provides a clear depiction of the Kuroshio pattern [37] bounded by the red box.
@@
To reveal the effect of learning Lagrangian trajectories, we visualize tracked particles in Figure 5. We observe that all the particles move from west to east, consistent with the Pacific circulation. Additionally, the tracked particles show distinct moving patterns, confirming their ability to represent complex dynamics. The movement of upper particles matches the sinuous trajectory of the Kuroshio current, demonstrating the capability of DeepLag to provide interpretable evidence for prediction results. Visualizing the tracking points in Lagrangian space instills confidence in the reliability and interpretability of the predictions made by our model, which can provide valuable and intuitive insights for real-world fluid dynamics.
@@
Table 4: Performance comparison on the 3D Smoke dataset. Relative L2 with relative promotion w.r.t. the second-best model is recorded.
@@
3D fluid prediction has been a long-standing challenge due to the tanglesome dynamics involved. Therefore, we generated this benchmark to describe a scenario in which smoke flows under the influence of buoyancy in a three-dimensional bounding box. For more details, please refer to Appendix C.3.
@@
In Figure 6, we compared the prediction results on the 3D Smoke dataset. DeepLag demonstrates superior performance in capturing both the convection and diffusion of smoke within the bounding box. In contrast, the predictions made by U-Net tend to average value across various surfaces, resulting in blurred details, which also indicates its deficiency in dynamics modeling. Similarly, LSM and FactFormer exhibit more pronounced errors, particularly around the smoke boundaries, where complex wave interactions often occur. By comparison, our model significantly reduces both the overall and cross-sectional error, excelling in the prediction of fine-grained, subtle flow patterns and maintaining high accuracy even in challenging regions with intricate dynamics.
Figure 6: Showcases comparison of the whole space (left part) and a cross-section (right part) on the 3D Smoke dataset. For better visualization, we present the absolute value of the prediction error and normalize the whole space error into $(0,0.12)$. As for the cross-section visualization, we choose the $xOy$ plane in the middle 3D fluid and normalize error maps to $(-0.5,0.5)$.